# MoC: Mixtures of Text Chunking Learners for Retrieval-Augmented Generation System

### v1

@@ 1 Introduction
Given the critical role of text chunking in RAG systems, optimizing this process has emerged as one of the key strategy to mitigate performance bottlenecks. Traditional text chunking methods, often based on rules or semantic similarity (Zhang et al., 2021; Langchain, 2023; Lyu et al., 2024), provide some structural segmentation but are inadequate in capturing subtle changes in logical relationships between sentences. The LumberChunker (Duarte et al., 2024) offers a novel solution by utilizing LLMs to receive a series of consecutive paragraphs and accurately identify where content begins to diverge. However, it demands a high level of instruction-following ability from LLMs, which incurs significant resource and time costs. Additionally, the effectiveness of current chunking strategies is often evaluated indirectly through downstream tasks, such as the QA accuracy in RAG systems, with a lack of independent metrics for evaluating the inherent rationality of the chunking process itself. These challenges give rise to two practical questions: This raises a practical question: How can we fully utilize the powerful reasoning capabilities of LLMs while accomplishing the text chunking task at a lower cost? And how to devise evaluation metrics that directly quantify the validity of text chunking?
@@ 2 Related Works
Text Chunking in RAG By expanding the input space of LLMs through introducing retrieved text chunks (Guu et al., 2020; Lewis et al., 2020), RAG significantly improves the performance of knowledge-intensive tasks (Ram et al., 2023). Text chunking allows information to be more concentrated, minimizing the interference of irrelevant information, enabling LLMs to focus more on the specific content of each text chunk and generate more precise responses (Yu et al., 2023; Besta et al., 2024; Su et al., 2024). LumberChunker (Duarte et al., 2024) iteratively harnesses LLMs to identify potential segmentation points within a continuous sequence of textual content, showing some potential for LLMs chunking. However, this method demands a profound capability of LLMs to follow instructions and entails substantial consumption when employing the Gemini model.
@@ 3.1 Deep Reflection on Chunking Strategies
As pointed out by Qu et al. (2024), semantic chunking has not shown a significant advantage in many experiments. This paper further explores this phenomenon and proposes two key metrics, "Boundary Clarity" and "Chunk Stickiness", to scientifically explain the limitations of semantic chunking and the effectiveness of LLM chunking. At the same time, it also provides independent evaluation indicators for the rationality of chunking itself.
@@ 3.1.1 Boundary Clarity (BC)
where $\text{ppl}(q)$ represents the perplexity of sentence sequence $q$, and $\text{ppl}(q|d)$ denotes the contrastive perplexity given the context $d$. Perplexity serves as a critical metric for evaluating the predictive accuracy of language models (LMs) on specific textual inputs, where lower perplexity values reveal superior model comprehension of the text, whereas higher values reflect greater uncertainty in semantic interpretation. When the semantic relationship between two text chunks is independent, $\text{ppl}(q|d)$ tends to be closer to $\text{ppl}(q)$, resulting in the BC metric approaching 1. Conversely, strong semantic interdependence drives the BC metric toward zero. Therefore, higher boundary clarity implies that chunks can be effectively separated, whereas a lower boundary clarity indicates blurred boundaries between chunks, which may potentially lead to information confusion and comprehension difficulties.
@@ 3.1.2 Chunk Stickiness (CS)
$\text{CS}(G)=-\sum_{i=1}^{n}\frac{d_{i}}{2m}\cdot\log_{2}{\frac{d_{i}}{2m}}$ (3)
where $G$ is the constructed semantic graph, $d_{i}$ represents the degree of node $i$, and $m$ denotes the total number of edges. This methodology constructs a complete graph, followed by redundancy reduction based on the inter-segment relationships.
@@ 4.4 Main Results
Upon validating the effectiveness of our proposed chunking experts, we proceeded to investigate the impact of various special characters on performance, and extended chunking within the MoC framework. As illustrated in Table 2, we design eight distinct special characters, each inducing varying degrees of performance fluctuation in the meta-chunker. Notably, all character configurations demonstrate measurable performance enhancements compared to baseline approaches, with $[Mask]$ and <.><.*>< . ∗ > exhibiting particularly remarkable efficacy. In our experiments, both the Meta-chunker-1.5B and the MoC framework employ $[Mask]$ as an ellipsis to replace the middle sections of text chunks, while maintaining consistent training data. The experimental results indicate that the chunking method based on the MoC architecture further enhances performance. Specifically, when handling complex long texts, MoC effectively differentiates the chunking granularity of various sections. Moreover, the time complexity of the MoC remains at the level of a single SLM, showcasing a commendable balance between computational efficiency and performance.
@@ 4.5 Exploring Chunking Based on Boundary Clarity and Chunk Stickiness
where $\text{sim}(q,d)$ represents the semantic similarity score between the text chunks $q$ and $d$. With this definition, the DS metric ranges from [0, 1], where 0 indicates perfect similarity and 1 indicates complete dissimilarity. The design of the DS metric is based on the following considerations: first, semantic similarity measures are typically employed to assess the degree of semantic proximity between two text segments. By converting this to the dissimilarity measure, we can more directly observe the semantic differences between chunks. Second, the linear transformation of DS preserves the monotonicity of the original similarity measure without losing any information. Figure 4(a) reveals the QA performance of RAG using different chunking strategies. It is important to note that, to ensure the validity of the evaluation, we maintained the same average text chunk length across all chunking methods.
Why Does Semantic Chunking Underperform? As illustrated in Figure 4(b), while semantic chunking scores are generally high, its performance in QA tasks is suboptimal. Moreover, there is no evident correlation between the scores of semantic dissimilarity and the efficacy of QA. This suggests that in the context of RAG, relying solely on semantic similarity between sentences is insufficient for accurately delineating the optimal boundaries of text chunks.
@@
Why Does LLM-Based Chunking Work? As shown in Table 3, the text chunks generated by LLMs exhibit superior boundary clarity, indicating the heightened ability to accurately identify semantic shifts and topic transitions, thereby mitigating the erroneous segmentation of related sentences. Concurrently, the LLM-based chunking produces text chunks with reduced chunk stickiness, signifying that the internal semantics of chunks are more tightly bound, while a greater degree of independence is maintained between chunks. This combination of well-defined boundaries and diminished stickiness contributes to enhanced retrieval efficiency and generation quality within RAG systems, ultimately leading to superior overall performance.
@@ 4.6 Hyper-parameter Sensitivity Analysis
Methods Complete Graph Incomplete Graph
@@
We conduct experiments on the decoding sampling hyperparameters of the meta-chunker within the MoC framework, with specific results presented in Table 2. Experimental data demonstrates that higher values of temperature and top-k sampling strategies introduce increased randomness, thereby exerting a certain impact on the chunking effect. Conversely, when these two hyperparameters are set to lower values, the model typically provides more stable and precise chunking, leading to a more significant performance improvement.
@@ A.1 Main Experimental Details
All language models utilized in this paper employ the chat or instruct versions where multiple versions exist, and are loaded in full precision (Float32). The vector database is constructed using Milvus, where the embedding model is bge-large-zh-v1.5. In experiments, we utilize a total of four benchmarks, and their specific configurations are detailed as follows:
@@ A.4 Another Perspective on Chunking Performance Comparison
A.5 Prompt utilized in Chunking

### v2

@@ 1 Introduction
Given the critical role of text chunking in RAG systems, optimizing this process has emerged as one of the key strategy to mitigate performance bottlenecks. Traditional text chunking methods, often based on rules or semantic similarity (Zhang et al., 2021; Langchain, 2023; Lyu et al., 2024), provide some structural segmentation but are inadequate in capturing subtle changes in logical relationships between sentences. The LumberChunker (Duarte et al., 2024) offers a novel solution by utilizing LLMs to receive a series of consecutive paragraphs and accurately identify where content begins to diverge. However, it demands a high level of instruction-following ability from LLMs, which incurs significant resource and time costs. Additionally, the effectiveness of current chunking strategies is often evaluated indirectly through downstream tasks, such as the QA accuracy in RAG systems, with a lack of independent metrics for evaluating the inherent rationality of the chunking process itself. These challenges give rise to two practical questions: How can we fully utilize the powerful reasoning capabilities of LLMs while accomplishing the text chunking task at a lower cost? And how to devise evaluation metrics that directly quantify the validity of text chunking?
@@ 2 Related Works
Text Chunking in RAG By expanding the input space of LLMs through introducing retrieved text chunks (Guu et al., 2020; Lewis et al., 2020), RAG significantly improves the performance of knowledge-intensive tasks (Ram et al., 2023). Text chunking allows information to be more concentrated, minimizing the interference of irrelevant information, enabling LLMs to focus more on the specific content of each text chunk and generate more precise responses (Yu et al., 2023; Besta et al., 2024; Su et al., 2024). LumberChunker (Duarte et al., 2024) iteratively harnesses LLMs to identify potential segmentation points within a continuous sequence of textual content, showing some potential for LLMs chunking. However, LumberChunker demands a profound capability of LLMs to follow instructions and entails substantial consumption when employing the Gemini model. Notably, access to official API interface of LLM involves cost implications, while common open-source models demonstrate limited generalization capabilities in the domain-specific task of text chunking.
@@ 3 Methodology
Addressing the current gap in independently evaluating chunking quality, this paper proposes two novel metrics: boundary clarity and chunk stickiness, as detailed in Section 3.1. Through a preliminary exploration of the chunking mechanism, we elucidate the necessity of leveraging LLMs for chunking tasks and introduce the MoC framework in Section 3.2. The framework employs a multi-granularity aware routing network with sparse activation to dynamically engage lightweight chunkers. Instead of generating complete text chunks, chunkers are guided to produce structured chunking rule lists, ensuring efficiency optimization without compromising chunking accuracy. A more comprehensive analysis is provided in Appendix A.6.
@@ 3.1 Deep Reflection on Chunking Strategies
As pointed out by Qu et al. (2024), semantic chunking has not shown a significant advantage in many experiments. This paper further explores this phenomenon and proposes two key metrics to scientifically explain the limitations of semantic chunking and the effectiveness of LLM chunking. At the same time, it also provides independent evaluation indicators for the rationality of chunking itself.
@@ 3.1.1 Boundary Clarity (BC)
where $\text{ppl}(q)$ represents the perplexity of sentence sequence $q$, and $\text{ppl}(q|d)$ denotes the contrastive perplexity given the text chunk $d$. Perplexity serves as a critical metric for evaluating the predictive accuracy of language models (LMs) on specific textual inputs, where lower perplexity values reveal superior model comprehension of the text, whereas higher values reflect greater uncertainty in semantic interpretation. When the semantic relationship between two text chunks is independent, $\text{ppl}(q|d)$ tends to be closer to $\text{ppl}(q)$, resulting in the BC metric approaching 1. Conversely, strong semantic interdependence drives the BC metric toward zero. Therefore, higher boundary clarity implies that chunks can be effectively separated, whereas a lower boundary clarity indicates blurred boundaries between chunks, which may potentially lead to information confusion and comprehension difficulties.
@@ 3.1.2 Chunk Stickiness (CS)
$\text{CS}(G)=-\sum_{i=1}^{n}\frac{h_{i}}{2m}\cdot\log_{2}{\frac{h_{i}}{2m}}$ (3)
where $G$ is the constructed semantic graph, $h_{i}$ represents the degree of node $i$, and $m$ denotes the total number of edges. This methodology constructs a complete graph, followed by redundancy reduction based on the inter-segment relationships.
@@ 4.4 Main Results
Upon validating the effectiveness of our proposed chunking experts, we proceeded to investigate the impact of various special characters on performance, and extended chunking within the MoC framework. As illustrated in Table 2, we design eight distinct special characters, each inducing varying degrees of performance fluctuation in the meta-chunker. Notably, all character configurations demonstrate measurable performance enhancements compared to baseline approaches, with $[Mask]$ and <.><.*>< . ∗ > exhibiting particularly remarkable efficacy. In our experiments, both the Meta-chunker-1.5B and the MoC framework employ $[Mask]$ as an ellipsis to replace the middle sections of text chunks, while maintaining consistent training data. The experimental results indicate that the chunking method based on the MoC architecture further enhances performance. Specifically, when handling complex long texts, MoC effectively differentiates the chunking granularity of various sections. Moreover, the time complexity of the MoC remains at the level of a single SLM, showcasing a commendable balance between computational efficiency and performance. Further details regarding analysis and experiments are outlined in Appendices A.4 and A.5.
@@ 4.5 Exploring Chunking Based on Boundary Clarity and Chunk Stickiness
where $\text{sim}(q,d)$ represents the semantic similarity score between the text chunks $q$ and $d$. With this definition, the DS metric ranges from [0, 1], where 0 indicates perfect similarity and 1 indicates complete dissimilarity. The design of the DS metric is based on the following considerations: first, semantic similarity measures are typically employed to assess the degree of semantic proximity between two text segments. By converting this to the dissimilarity measure, we can more directly observe the semantic differences between chunks. Second, the linear transformation of DS preserves the monotonicity of the original similarity measure without losing any information. The left side of Table 4 reveals the QA performance of RAG using different chunking strategies. It is important to note that, to ensure the validity of the evaluation, we maintained the same average text chunk length across all chunking methods.
Chunking Methods QA Performance Dissimilarity
BLEU-1 ROUGE-L Model-1 Model-2
Original 0.3515 0.4213 0.2731 0.2885
Llama_index 0.3620 0.4326 0.2836 0.3071
Semantic Chunking 0.3382 0.4131 0.4174 0.4162
Qwen2.5-14B 0.3650 0.4351 0.2955 0.3263
Table 4: QA performance across different chunking methods and the degree of dissimilarity between text chunks, with Model-1 and Model-2 denoting semantic similarity models bge-large-zh-v1.5 and all-MiniLM-L6-v2, Respectively.
Why Does Semantic Chunking Underperform? As illustrated on the right side of Table 4, while semantic chunking scores are generally high, its performance in QA tasks is suboptimal. Moreover, there is no evident correlation between the scores of semantic dissimilarity and the efficacy of QA. This suggests that in the context of RAG, relying solely on semantic similarity between sentences is insufficient for accurately delineating the optimal boundaries of text chunks. For further in-depth analyses and illustrations, refer to Appendix A.7 and Figure 7.
@@
Why Does LLM-Based Chunking Work? As shown in Table 3, text chunks generated by LLMs exhibit superior boundary clarity, indicating the heightened ability to accurately identify semantic shifts and topic transitions, thereby mitigating the erroneous segmentation of related sentences. Concurrently, the LLM-based chunking produces chunks with reduced chunk stickiness, signifying that the internal semantics of chunks are more tightly bound, while a greater degree of independence is maintained between chunks. We also validate that BC and CS exhibit a significant correlation with the QA performance, as detailed in Appendix A.7. Accordingly, this combination of well-defined boundaries and diminished stickiness contributes to enhanced retrieval efficiency and generation quality within RAG systems, ultimately leading to superior overall performance.
@@ 4.6 Hyper-parameter Sensitivity Analysis
Chunking Methods Complete Graph Incomplete Graph
@@
We conduct experiments on the decoding sampling hyperparameters of the meta-chunker within the MoC framework, with specific results presented in Figure 2. Experimental data demonstrates that higher values of temperature and top-k sampling strategies introduce increased randomness, thereby exerting a certain impact on the chunking effect. Conversely, when these two hyperparameters are set to lower values, the model typically provides more stable and precise chunking, leading to a more significant performance improvement.
@@ A.1 Main Experimental Details
All language models utilized in this paper employ the chat or instruct versions where multiple versions exist, and are loaded in full precision. The vector database is constructed using Milvus, where the embedding model is bge-large-zh-v1.5. In experiments, we utilize a total of four benchmarks, and their specific configurations are detailed as follows:
@@ A.3 Dataset Construction Process
Classification of Granularity Labels The division of granularity labels is strictly based on the statistical properties of the data distribution. Granularity labels are used to partition the original data into independent training sets, with a balanced amount of data corresponding to each label. Specifically, the total amount of data we processed is 20K. To ensure sufficient data under each label for adequate training of each chunker, we strive to maintain approximately 5K data points per label. In terms of independence, we ensure that each granularity label is relatively independent, meaning there is no overlap or significant correlation between labels, thus preventing confusion during the training process of chunkers.
Moreover, given the relative scarcity of datasets for text chunking, we have expanded our chunking dataset to 120K entries, with 60K entries each in Chinese and English, and have made it open-source for sharing within the academia. We are also continuously optimizing the dataset with the aim of further improving its quality. This will provide a more solid foundation for the division of granularity labels, ensuring that each chunker can be adequately trained with sufficient data support, thereby possessing greater robustness.
@@ A.4 Another Perspective on Chunking Performance Comparison
A.5 Exploration of Text Chunking Efficiency
To address the issue of complex and variable text chunking granularities in real-world scenarios, this paper proposes the multi-granularity chunking framework. This method schedules chunkers for different granularities through a dynamic routing mechanism. This sparse activation mechanism controls computational resource consumption at the level of a single chunker, enabling the overall system to optimize efficiency while maintaining accuracy.
  • This mechanism effectively resolves the granularity generalization dilemma faced by traditional single models, as shown in Figure 5: when training data covers a wide range of chunking granularities, SLMs with limited parameters struggle to capture multi-granularity feature correlations, while simply scaling up the model size leads to exponential growth in computational complexity.
  • The MoC framework decouples complex multi-granularity chunking tasks into multiple expert subtasks, allowing each chunker to focus solely on feature modeling within a specific granularity range.
  • It is worth emphasizing that our goal is not to require the model to generate every text chunk in its entirety, but rather to guide it in generating a structured list of chunking regular expressions. This approach offers dual advantages: on the one hand, it fully leverages the powerful reasoning capabilities of language models to ensure the reasonability of the chunking; on the other hand, by eliminating the need to generate the complete content of each text chunk, it saves generation time and improves overall processing efficiency.
To quantify the speed of chunking, we evaluate various chunking methods, which necessitate the utilization of models, using a single NVIDIA A800 GPU on the CRUD benchmark. The experimental results demonstrate that, although the MoC design incurs some additional computational overhead, we are able to effectively control the processing time while maintaining accuracy through the sparse activation mechanism and the use of 1.5B chunkers, as shown in Table 7. Furthermore, in RAG scenarios, chunking operations are usually used to construct a local vector knowledge base, which can be reused multiple times. Consequently, the emphasis is placed more on the precision of chunking in applications.
Chunking Methods Average Time Per Document (Unit: s)
Semantic Chunking 1.16
LumberChunker 3.23
Qwen2.5-14B 26.99
Meta-chunker-1.5B 3.69
Table 7: Efficiency comparison of different model-based chunking methods on the CRUD benchmark, with the unit defined as the average processing time in seconds per document.
A.6 Methodological Analysis
We firstly reorganize and briefly introduce the framework of the paper to convey our research ideas more clearly. This paper initially employs structured instruction design to explicitly guide the GPT-4o model [^2] https://openai.com/api/ in text chunking based on logical and semantic structures. In this process, we innovatively adopt a sliding window algorithm and a chunk buffering mechanism to handle long texts, ensuring information continuity. Simultaneously, to directly measure the quality of chunking, we propose two new metrics, boundary clarity and chunk stickiness, to strictly control the chunking results. Subsequently, when constructing the training dataset, we extract the first and last characters of text chunks as anchors, with the intermediate content replaced by special placeholders, thereby generating high-quality training samples. Our objective is to guide a SLM to generate a structured list of chunking regular expressions and then extract text chunks from the original text, resolving the conflict between LLM computational efficiency and chunking accuracy.
Figure 3: Score distribution of attention heads before fine-tuning.
Figure 4: Score distribution of attention heads after fine-tuning.
During the model training phase, the paper designs a multi-granularity aware router capable of dynamically selecting the most suitable chunker for documents requiring different chunk lengths, as depicted in Figure 5. Furthermore, we train a meta-chunker using a full fine-tuning strategy to generate structured chunking regular expressions. To address potential hallucination issues, we introduce an edit distance recovery algorithm, ensuring the accuracy of chunking. The entire process not only improves data quality and chunking precision but also maintains efficient computational capabilities through the sparse activation mechanism, replacing chunk generation with chunking rules generation, and utilizing smaller models in place of larger ones.
Chunking Methods CRUD (Single-hop) CRUD (Two-hop) DuReader WebCPM
BLEU-1 BLEU-Avg ROUGE-L BLEU-1 BLEU-Avg ROUGE-L F1 ROUGE-L
Original $0.3515_{-0.002}$ $0.2548_{-0.0014}$ $0.4213^{+0.0005}_{-0.0014}$ $0.2322^{+0.003}$ $0.1133^{+0.0023}$ $0.2613^{+0.0026}$ $0.2030^{+0.0029}$ $0.2642^{+0.0007}_{-0.004}$
Llama_index $0.3620_{-0.0109}$ $0.2682_{-0.0089}$ $0.4326_{-0.0115}$ $0.2315^{+0.0005}$ $0.1133^{+0.0016}$ $0.2585^{+0.0026}$ $0.2220_{-0.0076}$ $0.2630^{+0.0034}$
Semantic Chunking $0.3382^{+0.0007}_{-0.0012}$ $0.2462^{+0.0015}_{-0.0006}$ $0.4131^{+0.0005}_{-0.0012}$ $0.2223^{+0.0024}_{-0.0003}$ $0.1075_{-0.0012}$ $0.2507^{+0.0004}$ $0.2157_{-0.0061}$ $0.2691_{-0.0013}$
LumberChunker $0.3456_{-0.0052}$ $0.2542_{-0.0037}$ $0.4160_{-0.0037}$ $0.2204^{+0.0003}_{-0.0006}$ $0.1083^{+0.0008}_{-0.0004}$ $0.2521^{+0.0006}_{-0.001}$ $0.2178^{+0.0033}$ $0.2730^{+0.0024}$
[…]
Qwen2.5-72B $0.3721^{+0.0019}$ $0.2743^{+0.0015}$ $0.4405^{+0.0023}$ $0.2382_{-0.0053}$ $0.1185_{-0.004}$ $0.2677_{-0.0053}$ $0.2284^{+0.003}$ $0.2693^{+0.0017}$
[…]
Table 8: Confidence intervals of main results determined through three independent experiments.
On the other hand, within the current research context, traditional chunking methods exhibit several limitations. Specifically, existing chunking techniques typically require sentence segmentation prior to chunking sentences. This approach presents several evident issues:
  • Sentence boundaries in natural language are not always clear and unambiguous, particularly in texts containing extensive colloquial or non-standard usage, where automatic sentence segmentation may introduce errors.
  • Simple sentence segmentation methods fail to accurately process texts with complex structures, nested sentences, or long sentences. This can result in the omission or misinterpretation of crucial information.
  • Texts often contain structured information such as tables and code, and chunking by sentence may disrupt these structures, making it difficult to correctly understand and process the information.
Therefore, it is necessary to explore a novel chunking method to overcome the limitations of the segment-then-chunk approach. Driven by this motivation and necessity, we propose an end-to-end solution for text chunking using SLMs directly.
Notably, further discoveries are made when analyzing the attention score heatmaps of models within the MoC framework, as illustrated in Figures 3 and 4. Before fine-tuning, the attention heads assigned high scores to special tokens, but these scores decreased after fine-tuning, while the attention scores for characters at the beginning and end of text chunks increased. This indicates that before fine-tuning, the model’s attention allocation has not yet fully adapted to the specific task requirements. Some attention heads excessively focus on special tokens in the initial stage, which hinders the normal completion of the task. During the fine-tuning process, the model redistributes and adjusts its attention based on the data and objectives of the specific task. Attention heads that are initially overly focused on a certain token reduce their scores and shift their attention more towards information that is more relevant to the task. Consequently, through fine-tuning, the model’s understanding and mastery of the task improve, enabling it to more accurately identify and focus on the information that is truly important for task completion. This further demonstrates the effectiveness of our approach.
To ensure statistical robustness, we conduct three independent experiments for each method to determine the boundaries of confidence intervals, with some data showing a tendency towards one side. The specific experimental results are presented in Table 8.
Figure 5: Granularity distribution of text chunks generated by GPT-4o on the CRUD benchmark.
A.7 Relationship Between BC, CS, and the Performance of the RAG System
To more accurately demonstrate the correlation between BC, CS, and the performance of the RAG system, we add the calculation of Pearson correlation coefficients based on the existing experiments in the paper. The Pearson correlation coefficient is a statistical measure that assesses the degree of linear correlation between two variables, with a value range of [-1, 1]. A positive correlation coefficient indicates a positive linear relationship between the two variables; a negative correlation coefficient indicates a negative linear relationship; and a correlation coefficient of 0 indicates no correlation between the two variables. If the absolute value of the correlation coefficient is greater than 0.5, it suggests a relatively evident linear relationship between the variables.
We select the evaluation results of the Qwen2.5-7B, as shown in the table 9. After calculation, the correlation coefficients between BC, $\text{CS}_{c}$, $\text{CS}_{i}$, and ROUGE-L are found to be 0.8776, -0.7453, and -0.6663, respectively. This indicates that there is indeed a noticeable linear relationship between BC, CS, and the performance of the RAG system, further validating our discussion in the paper: BC represents boundary clarity, which is preferable when higher; $\text{CS}_{c}$ denotes chunk stickiness utilizing a complete graph, and $\text{CS}_{i}$ indicates chunk stickiness employing a incomplete graph, both of which are favorable when lower.
Chunking Methods ROUGE-L BC $\text{CS}_{c}$ $\text{CS}_{i}$
Original 0.4213 0.8049 2.421 1.898
Llama_index 0.4326 0.8455 2.250 1.483
Semantic Chunking 0.4131 0.8140 2.325 1.650
Qwen2.5-14B 0.4351 0.8641 2.125 1.438
Table 9: Performance of different chunking methods under Qwen2.5-7B.
A.8 Prompt utilized in Chunking